40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 4 10:54:12 2021, Paco, Anchal, Summary, LSC, POY11 measurement, tried to lock Green Yend laser 20210304_POY11_Spectrum_YARMLocked.pdf20210304_POY11_Spectrum_YARMLocked.tar.gz
    Reply  Thu Mar 4 11:59:25 2021, Paco, Anchal, Summary, LSC, Watchdog tripped, Optics damped back 
       Reply  Thu Mar 4 15:48:26 2021, Koji, Summary, PEM, Watchdog tripped, Optics damped back 
Message ID: 15861     Entry time: Thu Mar 4 10:54:12 2021     Reply to this: 15862
Author: Paco, Anchal 
Type: Summary 
Category: LSC 
Subject: POY11 measurement, tried to lock Green Yend laser 

[Paco, Anchal]

- First ran burtgooey as last time.

- Installed pyepics on base environment of donatella

ASS XARM:
- Clicked on ON in the drop down of "! More Scripts" below "! Scripts XARM" in C1ASS.adl
- Clicked on "Freeze Outputs" in the same menu after some time.
- Noticed that the sensing and output matrix of ASS on XARM and YARM look very different. The reason probably is because the YARM outputs have 4 TT1/2 P/Y dof instead of BS P/Y on the XARM. What are these TT1/2?

(Probably, unrelated but MC Unlocked and kept on trying to lock for about 10 minutes attaining the lock eventually.)

Locking XARM:
- From scripts/XARM we ran lockXarm.py from outside any conda environment using python command.
- Weirdly, we see that YARM is locked??? But XARM is not. Maybe this script is old.
- C1:LSC-TRY-OUTPUT went to around 0.75 (units unknown) while C1:LSC-TRX-OUTPUT is fluctuating around 0 only.

POY11 Spectrum measurement when YARM is locked:
- Created our own template as we couldn't find an existing one in users/Templates.
- Template file and data in Attachment 2.
- It is interesting to see most of the noise is in I quadrature with most noise in 10 to 100 Hz.
- Given the ARM is supposed to be much calmer than MC, this noise should be mostly due to the mode cleaner noise.
- We are not sure what units C1:LSC-POY11_I_ERR_DQ have, so Y scale is shown with out units.


Trying to lock Green YEND laser to YARM:
- We opened the Green Y shutter.
- We ensured that when temperature slider og green Y is moved up, the beatnote goes up.
- ARM was POY locked from previous step.
- Ran script scripts/YARM/Lock_ALS_YARM.py from outside any conda environment using python command.
- This locked green laser but unlocked the YARM POY.

Things moving around:
- Last step must have made all the suspension controls unstable.
- We see PRM and SRM QPDs moving a lot.
- Then we did burt restore to /opt/rtcds/caltech/c1/burt/autoburt/today/08:19/*.snap to go back to the state before we started changing things today.

[Paco left for vaccine appointment]

- However the unstable state didn't change from restore. I see a lot of movement in ITMX/Y. PRM and BS also now. Movement in WFS1 and MC2T as well.
 - I closed PSL shutter as well to hopefully disengage any loops that are still running unstably.
 - But at this point, it seems that the optics are just oscillating and need time to come back to rest. Hopefully we din't cause too much harm today :(.
 


My guess on what happened:

  • Us using the Lock_ALS_YARM.py probably created an unstable configuration in LSC matrix and was the start of the issue.
  • On seeing PRM fluctuate so much, we thought we should just burst restore everything. But that was a hammer to the problem.
  • This hammer probably changed the suspension position values suddenly causing an impulse to all the optics. So everything started oscillating.
  • Now MC WFS is waiting for MC to lock before it stablizes the mode cleaner. But MC autolocker is unable to lock because the optics are oscillating. Chicken-egg issue.
  • I'm not aware of how manually one can restore the state now. My only known guess is that if we wait for few hours, everything should calm back enough that MC can be locked and WFS servo can be switched on.
Attachment 1: 20210304_POY11_Spectrum_YARMLocked.pdf  34 kB  | Hide | Hide all
20210304_POY11_Spectrum_YARMLocked.pdf
Attachment 2: 20210304_POY11_Spectrum_YARMLocked.tar.gz  318 kB
ELOG V3.1.3-